Fermi National Laboratory


Accelerator Update for June 1, 2005 - June 3, 2005

Wednesday June 1
The day shift began with Operations monitoring store 4185 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with the MiniBooNE, NuMI, and the Meson MCenter experiments taking beam.

The experimenters at the Meson MTest beam line requested beam at 11:31 AM.

At 4:57 PM, a Booster RF station (BRF3) tripped off and caused the NuMI beam permit to drop. Everything reset okay.

MTest completed taking beam for the night at 8:37 PM.

At 10:03 PM, the NuMI horn tripped off. Operators conferred with a NuMI expert. They had the horn back on by 11:05 PM.

Thursday June 2
The midnight shift began with Operations monitoring store 4185 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations terminated store 4185 at 12:06 AM and put the Tevatron (TeV) into shot setup.

At 12:18 AM, the Booster notcher tripped off. An operator called in experts.

At 12:50 AM, the 400 MeV area chopper tripped off on an interlock trip. Operators couldn't reset the trip. The experts coming in to repair the notcher will investigate this problem.

Operations began preparing Linac for an access at 2:07 AM so the Booster experts could inspect the chopper.

A third Booster expert was called in at 2:17 AM to help with the Booster problems.

The Booster experts reported at 4:09 AM that the 400 MeV chopper and notcher were both back on line.

Operations established combination store 4187 at 6:38 AM with an initial luminosity of 103.5E30. Pbar resumed stacking and NuMI resumed taking beam.

At 8:49 AM, a Linac RF station (KRF1) tripped off and wouldn't reset. Operators investigated and found a problem with a LCW flow sensor. An expert had the flow sensor replaced by 9:17 AM.

Pbar resumed stacking at 10:03 AM.

Operations put the Recycler into shot setup at 11:41 AM to stash antiprotons.

The stash ended at 1:07 PM; Pbar resumed stacking and MCenter and MTest resumed taking beam.

At 7:40 PM, the Booster experienced a RF Anode Program Limit High trip. Operators were able to reset the trip, but these trips continued throughout the rest of the shift.

Operations disabled the MTest beam, as requested by the experimenters, at 7:23 PM.

The Recycler began stashing antiprotons at 8:25 PM. The stash was completed by 9:24 PM. Pbar resumed stacking and NuMI and MCenter resumed taking beam.

Friday June 3
The midnight shift began with Operations monitoring store 4187 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

At 3:36 AM, Operations reported that a NuMI power supply (V108E) tripped on a low limit. Operators reset the supply, but this happened repeatedly over the next hour. Operators swapped control cards with a good supply. Although this helped, there were still reference problems. At 5:05 AM, an operator called in a controls expert and Power supply experts. NuMI was off for the rest of the shift.

Operations terminated store 4187 at 8:06 AM.

The Plan for Friday and the Weekend
The plan for today is to prepare the TeV for access this morning to allow maintenance work. Operations will put the TeV into shot setup sometime after noon. The weekend plans are simple: stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/3/2005   email Fermilab